OpenStudy (anonymous):

find in terms of the constant a. lim h of 0 5(a+h)squared - (5a)squared ____________________________ h

OpenStudy (dumbcow):

First expand \[\rightarrow \frac{5(a^{2}+2ah+h^{2}) - 5a^{2}}{h}\] combine terms \[\rightarrow \frac{10ah +5h^{2}}{h}\] Factor out a h from numerator \[\rightarrow \frac{h(10a +5h)}{h} = 10a +5h\] Now take limit as h->0 \[\lim_{h \rightarrow 0} 10a+5h = 10a\]

myininaya (myininaya):

is that (5a)^2 or 5a^2?

OpenStudy (anonymous):

5a squared

myininaya (myininaya):

\[\lim_{h \rightarrow 0}\frac{5(a^2+2ah+h^2)-(5a)^2}{h}=\lim_{h \rightarrow 0}\frac{5a^2+10ah+5h^2-25a^2}{h}\] \[=\lim_{h \rightarrow 0}\frac{-20a^2+10ah+5h^2}{h} DNE\]

myininaya (myininaya):

is the square just on the a or is it on the whole thing because it looks like it is on the whole thing to me

OpenStudy (anonymous):

the whole thing

OpenStudy (dumbcow):

in that case, the first part should be ....(5(a+h))^2

myininaya (myininaya):

cow we aren't doing derivatives we are just finding limits

OpenStudy (dumbcow):

you will get 50a as the answer instead of 10 a ... looks like the definition of the derivative to me

myininaya (myininaya):

i believe anyways

OpenStudy (anonymous):

it reads 5 times a plus h squared minus 5a squared over h it says find in terms of the constant a

myininaya (myininaya):

\[\frac{5(a+h)^2-(5a)^2}{h} ?\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

or \[\frac{(5(a+h))^2-(5a)^2}{h}\]

OpenStudy (anonymous):

without () around 5a. NOT the second one you typed

OpenStudy (dumbcow):

well in that case as myininaya showed above...the limit does not exist or is undefined,

OpenStudy (dumbcow):

it doesn't make sense to me though...usually the coefficients would be the same, you just replace a with (a+h) .... f(a+h) - f(a) / h

OpenStudy (anonymous):

ok. THANK YOU GUYS! I have another one for you. I have the answer but Im not sure if its correct lim h of 0 the square root of (3+h) - the square root of 3 all over h. I got 1. is that wrong?

OpenStudy (dumbcow):

yeah the answer should be 1/(2sqrt(3))

OpenStudy (dumbcow):

when you multiplied by conjugate, make sure you do it on top and on bottom

OpenStudy (anonymous):

the answer DNE was incorrect

myininaya (myininaya):

if you want the right answer if you need to tell exactly what the problem is